OpenStudy (anonymous):

If A(0, 0) and B(2, 5), find AB. If necessary, submit your answer as a decimal to the nearest tenth. (Remember: AB means "the distance between points A and B.")

OpenStudy (kagıtucak):

AB=square root [(distance between x point)^2+(distance between y point^2)] so AB=root [(2-0)^2+(5-0)^2] AB=root (4+25) AB= root 29 AB= 5.3851

OpenStudy (anonymous):

thanks

OpenStudy (kagıtucak):

you're welcome